# Stabilization of active tissue deformation by a dynamic morphogen gradient

### v1

@@ Appendix F Deforming system with localized source
F.1 Discussion of the system size dynamics
@@
Starting with a gradient-extensile system with dimension $L_{x}=1/15$, we observe that the aspect ratio increases over time as expected (green curve in Figure 7A). However, the shear rate first increases and then decreases (green curve in Figure 7B), in contrast to the case without diffusion, where it monotonically decreases (Figure 6B). To understand this behavior, we first note that without any non-affine flows (i.e. $\bm{\mathrm{bar}{{\mathrm{v}}}}=\bm{0}$), both morphogen and source profiles keep their cosine profiles, $s(\bm{r},t)=\hat{s}[\cos{(2\pi x/L_{x})}+1]$ and $c(\bm{r},t)=\hat{c}(t)\cos{(2\pi x/L_{x})}+c_{\mathrm{offset}}(t)$, while only the morphogen amplitude $\hat{c}(t)$ and offset $c_{\mathrm{offset}}(t)$ vary over time. Without non-affine flows we can thus simplify the system dynamics to:
$\frac{\mathrm{d}\hat{c}}{\mathrm{d}t}$ $=-k_{\mathrm{d}}(\hat{c}-\hat{s})-4\pi^{2}D\frac{\hat{c}}{L_{x}^{2}}$ (81)
$\frac{\mathrm{d}L_{x}}{\mathrm{d}t}$ $=-\frac{2\pi^{2}\mathrm{sgn}(\alpha)\hat{c}^{2}}{L_{x}},$ (82)
where the offset $c_{\mathrm{offset}}(t)$ does not matter for the system size dynamics. We also noted that at each time point the morphogen amplitude $\hat{c}$ is relatively well approximated by its stationary value $\hat{c}_{\mathrm{s}}$ for given $L_{x}$ (Figure 9A). This corresponds to the adiabatic limit where the $c$ dynamics is fast as compared to the $L_{x}$ dynamics. This stationary value $\hat{c}_{\mathrm{s}}$ is given by Eq. (81) for $\mathrm{d}\hat{c}/\mathrm{d}t=0$, which implies:
$\hat{c}_{\mathrm{s}}=\left[1+\frac{4\pi^{2}D}{k_{\mathrm{d}}L_{x}^{2}}\right]^{-1}\hat{s}.$ (83)
Indeed, when numerically solving the system size dynamics given by Eq. (82) with the stationary solution for $\hat{c}_{\mathrm{s}}$, Eq. (83) instead of the actual $\hat{c}$ dynamics, Eq. (81), we get a similar behavior (Figure 9B). This behavior can be readily explained as initially $L_{x}$ is small, and thus the diffusion term in $\hat{c}$ dominates, such that $\hat{c}_{s}\sim L_{x}^{2}$. Inserting this into Eq. (82) gives a shear rate scaling of $(\mathrm{d}L_{x}/\mathrm{d}t)/L_{x}\sim L_{x}^{2}$, which leads for a gradient-extensile system ($\mathrm{sgn}(\alpha)=-1$) to an accelerating pure shear. Intuitively, this is because for small $L_{x}$, diffusion dominates, which generally lowers the amplitude of the morphogen field. In this limit, increasing $L_{x}$ decreases diffusion and thus increases the morphogen amplitude and, as a consequence, the shear rate. However, as soon as $L_{x}$ becomes large enough, the diffusion term in Eq. (83) can be neglected, and we have almost constant $\hat{c}$, such that the shear rate scales as $(\mathrm{d}L_{x}/\mathrm{d}t)/L_{x}\sim L_{x}^{-2}$. Hence, the shear decelerates over time due to the expansion of the morphogen field (subsection V.1).
For gradient-contractile systems, we observe a much more erratic system size behavior. The orange curves in Figure 7A,B show example trajectories for a system with initially $L_{x}=5$. In particular, the shear rate shows a transient, but pronounced peak before quickly switching to much slower dynamics (Figure 7B inset). This behavior can be quantitatively understood in terms of three distinct phases. In a first phase, $L_{x}$ is large, such that the diffusion term can be neglected in the morphogen dynamics, so that, according to Eq. (81), $\hat{c}$ keeps its initial value of approximately $\hat{s}$. Thus, according to Eq. (82), the shear rate initially scales as $(\mathrm{d}L_{x}/\mathrm{d}t)/L_{x}\sim-L_{x}^{-2}$, leading to a decrease in $L_{x}$ and an increase in the shear rate, as in the case without diffusion (subsection V.1). The second phase starts when $L_{x}$ becomes small enough for the diffusion term to dominate in the $\hat{c}$ dynamics, Eq. (81), such that $\mathrm{d}\hat{c}/\mathrm{d}t=-4\pi^{2}D\hat{c}/L_{x}^{2}$ (Figure 9E). Dividing Eq. (82) by $\mathrm{d}\hat{c}/\mathrm{d}t$, we get $\mathrm{d}L_{x}/\mathrm{d}\hat{c}=\mathrm{sgn}(\alpha)\hat{c}L_{x}/2D$, which implies that $L_{x}\sim\exp{[\mathrm{sgn}(\alpha)\hat{c}^{2}/4D]}$. This limit corresponds to the transient peak in the shear rate (Figure 9C,D). During this transient phase, $\hat{c}$ is rapidly decreasing, while for very small $\hat{c}$ the system dimension $L_{x}\sim\exp{[\mathrm{sgn}(\alpha)\hat{c}^{2}/4D]}$ varies only little as compared to $\hat{c}$. Thus, as $\hat{c}$ decreases, the diffusion term in Eq. (81) at some point becomes smaller than the source term $k_{\mathrm{d}}\hat{s}$, at which point this transient stops. During the subsequent third phase, the shear rate becomes small due to the small $\hat{c}$, which allows the morphogen dynamics to be close to its stationary state. But since $L_{x}$ is also small, the diffusion term still dominates in Eq. (83), and so the morphogen amplitude scales as $\hat{c}\sim L_{x}^{2}$, such that the shear rate scales as $(\mathrm{d}L_{x}/\mathrm{d}t)/L_{x}\sim L_{x}^{2}$, and thus slows down over time. Taken together, while the shear rate varies for both, gradient-extensile and gradient-contractile systems, its behavior is much more erratic in the gradient-contractile case.
@@ F.2 Affine diffusion-shear instability
$\omega\delta\tilde{c}(\mathrm{bar}q_{x})$ $=-k_{d}\big{(}1+\lambda^{2}q_{x}^{2}\big{)}\delta\tilde{c}(\mathrm{bar}q_{x})+2k_{d}\lambda^{2}q_{x}^{2}c_{0}(\mathrm{bar}q_{x})\frac{\delta\tilde{l}_{x}}{l_{x}}$ (84)
@@
Here, we also substituted the diffusion length scale $\lambda=\sqrt{D/k_{d}}$ and the lab-frame wave vector $q_{x}=\mathrm{bar}q_{x}/l_{x}$. Iff this linear system of equations has a positive eigenvalue $\omega$, then there is an instability.
@@ F.2.1 Cosine source profile
$\omega\delta\tilde{c}$ $=-k_{d}\big{(}1+\lambda^{2}q_{0}^{2}\big{)}\delta\tilde{c}+2k_{d}\lambda^{2}q_{0}^{2}c_{0}\frac{\delta\tilde{l}_{x}}{l_{x}}$ (86)
@@
where $q_{0}(t):=\mathrm{bar}q_{0}/l_{x}(t)=2\pi/L_{x}(t)$. There are two eigenvalues, $\omega_{1}$ and $\omega_{2}$ with $\omega_{1}\leq\omega_{2}$, and the larger eigenvalue is given by: The larger of the two eigenvalues is:
@@
where $\mathcal{T}:=-k_{d}(1+\lambda^{2}q_{0}^{2})+2\mathrm{sgn}(\alpha)q_{0}^{2}|c_{0}|^{2}$ and $\mathcal{D}:=-2\mathrm{sgn}(\alpha)k_{d}q_{0}^{2}|c_{0}|^{2}(1-3\lambda^{2}q_{0}^{2})$ are trace and determinant, respectively, of the matrix on the right-hand side of Eqs. (86) and (87).
In the gradient-extensile case, $\mathrm{sgn}(\alpha)=-1$, we have that $\omega_{1}+\omega_{2}=\mathcal{T}<0$. Thus, at least one eigenvalue is negative. The product of both eigenvalues is given by $\omega_{1}\omega_{2}=\mathcal{D}=2k_{d}q_{0}^{2}|c_{0}|^{2}(1-3\lambda^{2}q_{0}^{2})$. Hence, there is a positive eigenvalue iff $13\lambda^{2}q_{0}^{2}>$ 1, i.e. iff $2\pi\lambda\sqrt{3}>L_{x}$.
@@ F.2.2 Arbitrary source profile
$\delta\tilde{c}(\mathrm{bar}q_{x})=\frac{2\lambda^{2}q_{x}^{2}c_{0}(\mathrm{bar}q_{x})}{1+\omega/k_{d}+\lambda^{2}q_{x}^{2}}\,\frac{\delta\tilde{l}_{x}}{l_{x}}.$ (89)
@@
$\begin{split}\omega&=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{4\lambda^{2}q_{x}^{2}}{1+\omega/k_{d}+\lambda^{2}q_{x}^{2}}}-\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}}\\ &=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\omega/k_{d}}{1+\omega/k_{d}+\lambda^{2}q_{x}^{2}}}.\end{split}$ (90)
To solving this self-consistent equation for $\omega$, we look for the zeros of the function
$f_{\lambda}(\omega):=-\omega+\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\omega/k_{d}}{1+\omega/k_{d}+\lambda^{2}q_{x}^{2}}}.$ (91)
Specifically, we now show that iff $f_{\lambda}(0)>0$, there exits a positive solution in $\omega$, i.e. an $\omega>0$ for which $f_{\lambda}(\omega)=0$. This means that there is an affine diffusion-shear instability iff $f_{\lambda}(0)>0$. To show this, we first assume that there is an $\omega>0$ for which $f_{\lambda}(\omega)=0$. Then, we have:
$\begin{split}0&<\omega\\ &\quad=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\omega/k_{d}}{1+\omega/k_{d}+\lambda^{2}q_{x}^{2}}}\\ &\quad<\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1}{1+\lambda^{2}q_{x}^{2}}}\\ &\qquad=f_{\lambda}(0).\end{split}$ (92)
@@
Second, we assume that $f_{\lambda}(0)>0$. To show that there is a positive zero of $f_{\lambda}$, we need to show that there is a positive $\hat{\omega}>0$ for which $f_{\lambda}(\hat{\omega})<0$. Then, the Zwischenwertsatz implies that there is an $\omega$ with $0<\omega<\hat{\omega}$ such that $f_{\lambda}(\omega)=0$. Indeed, for $\hat{\omega}=3\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}}$, we find that:
$\begin{split}f_{\lambda}(\hat{\omega})&=-3\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}}\\ &\qquad\qquad\qquad+\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\hat{\omega}/k_{d}}{1+\hat{\omega}/k_{d}+\lambda^{2}q_{x}^{2}}}\\ &=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{-2(1+\hat{\omega}/k_{d})}{1+\hat{\omega}/k_{d}+\lambda^{2}q_{x}^{2}}}\\ &<0.\end{split}$ (93)
@@
Choosing for $s_{0}$ the von Mises profile in Eq. (20) with finite width $\mathrm{{w}}$, we numerically obtain $f_{\lambda}(0)>0$ iff $\lambda>g\mathrm{{w}}$ with $g\approx 0.639$ in the limit where $\mathrm{{w}}\ll L_{x}$ (i.e. where the von Mises profile becomes a Gaussian with standard deviation $\mathrm{{w}}$). However, the precise value of $g$ depends on the shape of the profile $s_{0}$. For a rectangular profile of $s_{0}$, we numerically even found that the instability always occurs.

### v2

@@ Appendix F Deforming system with localized source
F.1 System dimension dynamics
The dynamics of system dimension $L_{x}$, and thus box shear $l_{x}$, is given by Eq. (7).
We first discuss the case without diffusion, $D=0$, where source and signaling fields, $s(\mathrm{bar}{x})$ and $c(\mathrm{bar}{x})$, are stationary in co-deforming coordinates. In this case, the shear rate is essentially given by the integral over $(\partial_{x}c_{0})^{2}$. Because the signaling profile, $c_{0}$, is stationary in co-deforming coordinates, we have that for varying system size, $(\partial_{x}c_{0})^{2}\sim 1/l_{x}^{2}(\mathrm{bar}\partial_{x}c_{0})^{2}\sim 1/l_{x}^{2}$. So, using dimensionless units (subsection II.3), we can write
$\frac{1}{l_{x}}\,\frac{\mathrm{d}l_{x}}{\mathrm{d}t}=-\frac{\mathrm{sgn}(\alpha)}{L_{x}^{2}(0)l_{x}^{2}}.$ (77)
The solution of this is:
$l_{x}(t)=\sqrt{1-\frac{2\mathrm{sgn}(\alpha)t}{L_{x}(0)^{2}}}.$ (78)
Thus, gradient-extensile systems ($\alpha<0$) deform more and more slowly over time due to the widening of the signaling profile (green curves in Figure 7A). Meanwhile, gradient-contractile systems ($\alpha>0$) deform faster and faster over time due to the compression of the signaling profile, until the deformation rate diverges at some critical time point $t_{\mathrm{crit}}=L_{x}^{2}(0)/2$ (orange curves in Figure 7A).
Figure 7: System size dynamics of a freely deforming system in the limit $\mathrm{{w}}\rightarrow\infty$, where the source has a cosine shape. Here, $k_{\mathrm{d}}=1$. (A) Aspect ratios (left) and shear rates (right) over time for the gradient-extensile (green) and contractile (orange) systems, starting at initial aspect ratios of $L_{x}(0)/L_{y}(0)=1/225$ and $L_{x}(0)/L_{y}(0)=25$, respectively. Here the signaling molecule is non-diffusive, $D=0$. (B) The same data as panel A shown for a system with diffusive signaling molecule with $D=10^{-2}$. Here, the gradient-extensile (green) and gradient-contractile (orange) systems start from $L_{x}(0)/L_{y}(0)=0.04$ and $L_{x}(0)/L_{y}(0)=25$, respectively. Inset: zoomed around the sharp drop in the gradient-contractile case.
This box dimension dynamics is changed when allowing for finite diffusion, $D>0$ (Figure 7B). Specifically, the shear rate of gradient-extensile systems first increases, before it decreases like in the case without diffusion (compare green curves in Figure 7A,B, right panels). In gradient contractile systems, the shear rate first increases, like in the case without diffusion before it rather abruptly drops (compare orange curves in Figure 7A,B right panels and Figure 7B inset).
@@ F.1 System dimension dynamics
To better understand these effects of diffusion, we first note that without any non-affine flows (i.e. $\bm{\mathrm{bar}{{\mathrm{v}}}}=\bm{0}$), both signaling and source profiles keep their cosine profiles, $s(\bm{r},t)=\hat{s}[\cos{(2\pi x/L_{x})}+1]$ and $c(\bm{r},t)=\hat{c}(t)\cos{(2\pi x/L_{x})}+c_{\mathrm{offset}}(t)$, while only the signaling amplitude $\hat{c}(t)$ and offset $c_{\mathrm{offset}}(t)$ vary over time. We can thus simplify the dynamics to:
$\frac{\mathrm{d}\hat{c}}{\mathrm{d}t}$ $=-k_{\mathrm{d}}(\hat{c}-\hat{s})-4\pi^{2}D\frac{\hat{c}}{L_{x}^{2}(0)l_{x}^{2}}$ (79)
$\frac{1}{l_{x}}\,\frac{\mathrm{d}l_{x}}{\mathrm{d}t}$ $=-\frac{2\pi^{2}\mathrm{sgn}(\alpha)\hat{c}^{2}}{L_{x}^{2}(0)l_{x}^{2}},$ (80)
where $\hat{s}$ is constant since $\bm{\mathrm{bar}{{\mathrm{v}}}}=\bm{0}$, and the offset $c_{\mathrm{offset}}(t)$ does not matter for the box dimension dynamics. Eq. (80) implies that the shear rate scales as $\sim\hat{c}^{2}/l_{x}^{2}$. Further, the diffusion term in Eq. (79) can be neglected as long as $l_{x}\gtrsim\sqrt{D/k_{\mathrm{d}}}/L_{x}(0)$, and consequentially $\hat{c}\simeq\hat{s}$. In other words, the box shear rate for large $l_{x}$ is the same as in the case without diffusion, where the shear rate scales as $\sim l_{x}^{-2}$, for both gradient-extensile and gradient-contractile systems, consistent with Figure 7A,B. Conversely, taking Eq. (79) to the adiabatic limit of large diffusion $D$ and/or small box shear $l_{x}$, we have $\hat{c}\sim(k_{\mathrm{d}}/D)L_{x}^{2}(0)l_{x}^{2}\hat{s}\sim l_{x}^{2}$. Thus, in the limit of small box shear $l_{x}$, the shear rate scales as $\sim\hat{c}^{2}/l_{x}^{2}\sim l_{x}^{2}$, i.e. it increases with $l_{x}$, again consistent with Figure 7A,B.
@@ F.2 Affine diffusion-shear instability
$\omega\delta\tilde{c}(\mathrm{bar}q_{x})$ $=-k_{\mathrm{d}}\big{(}1+\lambda^{2}q_{x}^{2}\big{)}\delta\tilde{c}(\mathrm{bar}q_{x})+2k_{\mathrm{d}}\lambda^{2}q_{x}^{2}c_{0}(\mathrm{bar}q_{x})\frac{\delta\tilde{l}_{x}}{l_{x}}$ (81)
@@
Here, we also substituted the diffusion length scale $\lambda=\sqrt{D/k_{\mathrm{d}}}$ and the lab-frame wave vector $q_{x}=\mathrm{bar}q_{x}/l_{x}$. Iff this linear system of equations has a positive eigenvalue $\omega$, then there is an instability.
@@ F.2.1 Cosine source profile
$\omega\delta\tilde{c}$ $=-k_{\mathrm{d}}\big{(}1+\lambda^{2}q_{0}^{2}\big{)}\delta\tilde{c}+2k_{\mathrm{d}}\lambda^{2}q_{0}^{2}c_{0}\frac{\delta\tilde{l}_{x}}{l_{x}}$ (83)
@@
where $q_{0}(t):=\mathrm{bar}q_{0}/l_{x}(t)=2\pi/L_{x}(t)$. There are two eigenvalues, $\omega_{1}$ and $\omega_{2}$ with $\omega_{1}\leq\omega_{2}$, and the largest is given by:
@@
where $\mathcal{T}:=-k_{\mathrm{d}}(1+\lambda^{2}q_{0}^{2})+2\mathrm{sgn}(\alpha)q_{0}^{2}|c_{0}|^{2}$ and $\mathcal{D}:=-2\mathrm{sgn}(\alpha)k_{\mathrm{d}}q_{0}^{2}|c_{0}|^{2}(1-3\lambda^{2}q_{0}^{2})$ are trace and determinant, respectively, of the matrix on the right-hand side of Eqs. (83) and (84).
In the gradient-extensile case, $\mathrm{sgn}(\alpha)=-1$, we have that $\omega_{1}+\omega_{2}=\mathcal{T}<0$. Thus, at least one eigenvalue is negative. The product of both eigenvalues is given by $\omega_{1}\omega_{2}=\mathcal{D}=2k_{\mathrm{d}}q_{0}^{2}|c_{0}|^{2}(1-3\lambda^{2}q_{0}^{2})$. Hence, there is a positive eigenvalue iff $3\lambda^{2}q_{0}^{2}>1$, i.e. iff $2\pi\lambda\sqrt{3}>L_{x}$.
@@ F.2.2 Arbitrary source profile
$\delta\tilde{c}(\mathrm{bar}q_{x})=\frac{2\lambda^{2}q_{x}^{2}c_{0}(\mathrm{bar}q_{x})}{1+\omega/k_{\mathrm{d}}+\lambda^{2}q_{x}^{2}}\,\frac{\delta\tilde{l}_{x}}{l_{x}}.$ (86)
@@
$\begin{split}\omega&=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{4\lambda^{2}q_{x}^{2}}{1+\omega/k_{\mathrm{d}}+\lambda^{2}q_{x}^{2}}}-\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}}\\ &=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\omega/k_{\mathrm{d}}}{1+\omega/k_{\mathrm{d}}+\lambda^{2}q_{x}^{2}}}.\end{split}$ (87)
To solve this self-consistent equation for $\omega$, we look for the zeros of the function
$f_{\lambda}(\omega):=-\omega+\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\omega/k_{\mathrm{d}}}{1+\omega/k_{\mathrm{d}}+\lambda^{2}q_{x}^{2}}}.$ (88)
Specifically, we now show that iff $f_{\lambda}(0)>0$, there exists a positive solution in $\omega$, i.e. an $\omega>0$ for which $f_{\lambda}(\omega)=0$. This means that there is an affine diffusion-shear instability iff $f_{\lambda}(0)>0$. To show this, we first assume that there is an $\omega>0$ for which $f_{\lambda}(\omega)=0$. Then, we have:
$\begin{split}0&<\omega\\ &\quad=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\omega/k_{\mathrm{d}}}{1+\omega/k_{\mathrm{d}}+\lambda^{2}q_{x}^{2}}}\\ &\quad<\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1}{1+\lambda^{2}q_{x}^{2}}}\\ &\qquad=f_{\lambda}(0).\end{split}$ (89)
@@
Second, we assume that $f_{\lambda}(0)>0$. To show that there is a positive zero of $f_{\lambda}$, we need to show that there is a positive $\hat{\omega}>0$ for which $f_{\lambda}(\hat{\omega})<0$. Then, the Intermediate Value Theorem implies that there is an $\omega$ with $0<\omega<\hat{\omega}$ such that $f_{\lambda}(\omega)=0$. Indeed, for $\hat{\omega}=3\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}}$, we find that:
$\begin{split}f_{\lambda}(\hat{\omega})&=-3\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}}\\ &\qquad\qquad\qquad+\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{3\lambda^{2}q_{x}^{2}-1-\hat{\omega}/k_{\mathrm{d}}}{1+\hat{\omega}/k_{\mathrm{d}}+\lambda^{2}q_{x}^{2}}}\\ &=\sum_{\mathrm{bar}q_{x}}{q_{x}^{2}|c_{0}(\mathrm{bar}q_{x})|^{2}\frac{-2(1+\hat{\omega}/k_{\mathrm{d}})}{1+\hat{\omega}/k_{\mathrm{d}}+\lambda^{2}q_{x}^{2}}}\\ &<0.\end{split}$ (90)
@@
Choosing for $s_{0}$ the von Mises profile in Eq. (21) with finite width $\mathrm{{w}}$, we numerically obtain $f_{\lambda}(0)>0$ iff $\lambda>g\mathrm{{w}}$ with $g\approx 0.639$ in the limit where $\mathrm{{w}}\ll L_{x}$ (i.e. where the von Mises profile becomes a Gaussian with standard deviation $\mathrm{{w}}$). However, the precise value of $g$ depends on the shape of the profile $s_{0}$. For a rectangular profile of $s_{0}$, we numerically even find that the instability always occurs.
Appendix G Non-linearity in the active stress
We consider the system of equations, Eqs. (1)–(6), where we introduce a non-linearity into the active stress, Eq. (6):
$\tilde{\sigma}^{a}_{ij}=\alpha\,h\big{(}[\partial_{k}c]^{2}\big{)}\left[(\partial_{i}c)(\partial_{j}c)-\frac{1}{2}(\partial_{k}c)^{2}\delta_{ij}\right],$ (92)
where $z\mapsto h(z)$ is some real function expressing the nonlinearity. In the following, we assume that $h(z)>0$ for any real $z$, and that the function $z\mapsto zh(z)$ is monotonously increasing, i.e. the magnitude of the active stress increases monotonously with the magnitude of scalar field gradient.
Then, the dimensionless, linearized dynamics become:
$\frac{\mathrm{d}\delta l_{x}(t)}{\mathrm{d}t}$ $=-2\mathrm{sgn}(\alpha)\iint{\Big{[}h(z)+zh^{\prime}(z)\Big{]}\frac{c_{0}^{\prime}}{l_{x0}}\,\delta c^{\prime}\,\mathrm{d}\mathrm{bar}x\mathrm{d}\mathrm{bar}y}$
$\qquad+\mathrm{sgn}(\alpha)\big{(}I_{0}+2\Delta I\big{)}\,\delta l_{x},$ (95)
where $\delta c^{\prime}:=\mathrm{bar}\partial_{x}\delta c$, and
$I_{0}$ $:=\iint{zh(z)\,\mathrm{d}\mathrm{bar}x\mathrm{d}\mathrm{bar}y}$ (96)
$\Delta I$ $:=\iint{z^{2}h^{\prime}(z)\,\mathrm{d}\mathrm{bar}x\mathrm{d}\mathrm{bar}y}$ (97)
with $z(\bm{\mathrm{bar}{r}}):=(c_{0}^{\prime}(\bm{\mathrm{bar}{x}})/l_{x0})^{2}$.
We further focus on the gradient-extensile case, $\mathrm{sgn}(\alpha)=-1$, without non-affine flows, $\bm{\mathrm{bar}{\mathrm{v}}}=\bm{0}$, and, for simplicity, we consider the limit $\mathrm{{w}}\rightarrow\infty$, such that $s_{0}(\mathrm{bar}{x})\sim\cos{(\mathrm{bar}{q}_{0}\mathrm{bar}{x})}$, and thus also $c_{0}(\mathrm{bar}{x})=\hat{c}_{0}\cos{(\mathrm{bar}{q}_{0}\mathrm{bar}{x})}$ with $\mathrm{bar}{q}_{0}=2\pi/L_{x}(0)$ and some constant prefactor $\hat{c}_{0}$. As a consequence, based on Eq. (94), all Fourier modes of $\delta c$ decay except for $\delta c=\delta\hat{c}\cos{(\mathrm{bar}{q}_{0}\mathrm{bar}{x})}$. Hence, $\delta c(\mathrm{bar}{x})=(\delta\hat{c}/\hat{c}_{0})c_{0}(\mathrm{bar}{x})$. As a consequence, Eqs. (94) and (95) simplify to:
$\frac{\mathrm{d}\delta\hat{c}}{\mathrm{d}t}$ $=-(k_{\mathrm{d}}+Dq_{0}^{2})\delta\hat{c}+2Dq_{0}^{2}\hat{c}_{0}\frac{\delta l_{x}}{l_{x0}},$ (98)
$\frac{\mathrm{d}\delta l_{x}(t)}{\mathrm{d}t}$ $=2\big{(}I_{0}+\Delta I\big{)}l_{x0}\frac{\delta\hat{c}}{\hat{c}_{0}}-\big{(}I_{0}+2\Delta I\big{)}\,\delta l_{x}.$ (99)
This is a linear system of differential equations with constant coefficients. It is linearly stable if both eigenvalues, $\omega_{1}$ and $\omega_{2}$, of the coefficient matrix are negative. This implies that the trace of the matrix needs to be negative: $0>\omega_{1}+\omega_{2}=-k_{\mathrm{d}}-Dq_{0}^{2}-I_{0}-2\Delta I$, and that the determinant of the matrix needs to be positive: $0<\omega_{1}\omega_{2}=(k_{\mathrm{d}}+Dq_{0}^{2})(I_{0}+2\Delta I)-4Dq_{0}^{2}(I_{0}+\Delta I)=k_{\mathrm{d}}(I_{0}+2\Delta I)-Dq_{0}^{2}(3I_{0}+2\Delta I)$. For the system to be stable also for large diffusion, i.e. for $k_{\mathrm{d}}\ll Dq_{0}^{2}$, the latter condition implies $3I_{0}-2\Delta I>0$. However, we have assumed that $h(z)>0$, implying $I_{0}>0$, and that $z\mapsto zh(z)$ increases monotonically, implying that $h(z)+zh^{\prime}(z)\geq 0$ and thus $I_{0}+\Delta I\geq 0$. This contradicts determinant condition for a stable system. Taken together, this means that the diffusion-shear instability arises for large diffusion for any non-linearity $h$ in the active stress. We expect this result to generalize also to finite source widths $\mathrm{{w}}$; higher-order Fourier modes in $c_{0}$ would increase prefactor of the first term in Eq. (99), and thus make the system more unstable.